OpenStudy (anonymous):

Medal The dosage of a certain infant antibiotic prescribed by a pediatrician is determined by the weight of the infant in kilograms. If an infant who weighs 4.5 kilograms is given a shot containing 180 milligrams of this antibiotic, find the dosage rate for this medication in milligrams per kilogram (mg/kg). ? mg/kg

OpenStudy (anonymous):

divide

OpenStudy (tkhunny):

kilograms are mass, not weight. Whatever...

OpenStudy (anonymous):

per means divided by

OpenStudy (texaschic101):

180 mg to 4.5 kg make a proportion 180/4.5 = x/1 --- 180 mg to 4.5 kg = 1 mg to x kg cross multiply 4.5x = 180 x = 180/4.5 x = 40 mg per 1 kg...or 40/1

OpenStudy (texaschic101):

oops..explained proportion wrong 180/4.5 = x/1 --- 180 mg to 4.5 kg = x mg to 1 kg

OpenStudy (anonymous):

so its ok 40